# Attend and Enrich: Enhanced Visual Prompt for Zero-Shot Learning

### v2

@@ Introduction
As shown in Table 2, we evaluate key components in AENet, i.e., the learnable prompt $P$, semantic-related details $Z$ for the semantic-enhanced prompt, and concept-aware attention for the concept-harmonized token generation. When the original learnable prompt is removed, the model can be regarded as the baseline that directly uses the visual features extracted from vanilla ViT and projects them into semantic space for category inference. AENet outperforms the baseline with large $acc$/$H$ margins of 20.1%/6.7%, 5.9%/6.2%, and 4.5%/2.8% on CUB, SUN, and AwA2 datasets, respectively. The performance of AENet degrades when residual details are removed, as evidenced by decreases in both ZSL and GZSL across multiple datasets: specifically, a reduction in $acc$/$H$ of 4.5%/2.2% on CUB, 2.4%/4.8% on SUN, and 2.0%/0.9% on AwA2, compared to the full model implementation. Additionally, concept-aware attention provides concept-harmonized visual and semantic representations, resulting in $acc$/$H$ improvements of 1.5%/0.9%, 1.6%/2.4%, 1.5%/1.4% on CUB, SUN, and AwA2 datasets, respectively.
Figure 4: Effect of loss hyper-parameters on (a) CUB (b) SUN and (c) AwA2 datasets.
@@
In this work, we combine the attribute consistency loss $\mathcal{L}_{cons}$ and $\mathcal{L}_{deb}$ and debiasing with the balance hyper-parameters of $\lambda_{cons}$ and $\lambda_{deb}$ respectively to train the model. Here, we study the impact of $\lambda_{cons}$ and $\lambda_{deb}$ as shown in Figure 4. As $\lambda_{cons}$ rises from 0.0 to 2.0, i.e., the attribute-supervision $\mathcal{L}_{cons}$ is introduced into AENet for the semantic prediction, $H$ increases in all datasets. The best $H$ is obtained when $\lambda_{cons}=1.0$. This demonstrates the effectiveness of semantic consistency in predicting semantically relevant details, closely aligning with category prototypes to focus on attribute-specific features. When $\lambda_{cons}>1.0$, $H$ starts to drop. Thus, we set $\lambda_{cons}=1.0$ for optimal results. Additionally, when we gradually increase the value of $\lambda_{deb}$, more attention is paid to pursuing the consistent distribution between seen and unseen predictions, resulting in improved unseen accuracy $U$ and ultimately achieving better $H$ performance.
@@
To intuitively validate the capability of our AENet for capturing enhanced semantic information necessary for unseen classes, we input an unseen image into AENet to predict the attribute score ($\left\langle{\mathcal{M}(f(x)),{a_{\hat{y}(\hat{y}\in\mathcal{Y}^{u})}}}\right\rangle$). As shown in Figure 5, we can observe that AENet demonstrates strong performance in predicting attributes across various animal species in the AwA2 dataset. For example, for the horse, AENet closely matches the ground truth (GT) on “brown”, “hooves”, and “domestic”, with slight discrepancies in “chewteeth”, “smelly”, and “oldworld”. The rat predictions are remarkably accurate across all attributes, showcasing AENet’s precision for this species. These results illustrate AENet’s robust capability in capturing enhanced semantic information across diverse animal species, crucial for generalizing to unseen classes in zero-shot learning tasks.
@@
As shown in Figure 6, we present the t-SNE visualization of visual features for seen and unseen classes on AwA2. Compared to our full AENet, visual features extracted from the AENet w/o $P$ lack distinctiveness within certain classes. This observation intuitively suggests that the prompt plays a crucial role in enabling the application of pre-trained ViT to downstream ZSL tasks by generating high-quality features for seen classes. Furthermore, compared to AENet w/o $Z$, the visual features learned by our full AENet showcase desirable distinguishability with higher inter-class discrepancy and clearer decision boundaries.
@@ Conclusion
This paper introduces AENet, which leverages prompt learning and addresses the challenge of over-emphasizing primary visual features from seen domains. AENet comprises two key steps. Firstly, the concept-harmonized tokens are explored for visual and attribute modalities based on the modal-sharing token with their consistent concepts. Secondly, by investigating the semantic-related details, the VRRU is proposed to distill the semantic-enhanced prompt integrated with primary visual features to attend to semantic-related information for visual enrichment. Consequently, AENet can achieve superior visual-semantic alignment for generalizing to unseen classes. Extensive experiments across three benchmark datasets show the superiority of our AENet.

### v3

@@ Introduction
As shown in Table 2, we evaluate key components in AENet, i.e., the learnable prompt $P$, semantic-related details $Z$ for the semantic-enhanced prompt, and concept-aware attention for the concept-harmonized token generation. When the original learnable prompt is removed, the model can be regarded as the baseline that directly uses the visual features extracted from vanilla ViT and projects them into semantic space for category inference. AENet outperforms the baseline with large $acc$/$H$ margins of 20.1%/6.7%, 5.9%/6.2%, and 4.5%/2.8% on CUB, SUN, and AwA2 datasets, respectively. The performance of AENet degrades when residual details are removed, as evidenced by decreases in both ZSL and GZSL across multiple datasets: specifically, a reduction in $acc$/$H$ of 4.5%/2.2% on CUB, 2.4%/4.8% on SUN, and 2.0%/0.9% on AwA2, compared to the full model implementation. Additionally, concept-aware attention provides concept-harmonized visual and semantic representations, resulting in $acc$/$H$ improvements of 1.5%/0.9%, 1.6%/2.4%, 1.5%/1.4% on CUB, SUN, and AwA2 datasets, respectively. Additionally, we implement variants by replacing ZLinear in Eq. (4) with MLP and a linear layer, and the skip connection in Eq. (6) with a gated skip. As demonstrated in Table 3, these more complex alternatives do not offer performance improvements as the harmonized tokens are concise and informative. Moreover, ZLinear consistently achieves better performance compared to the linear.
Figure 4: Effect of loss hyper-parameters.
@@
In this work, we combine the attribute consistency loss $\mathcal{L}_{cons}$ and $\mathcal{L}_{deb}$ and debiasing with the balance hyper-parameters of $\lambda_{cons}$ and $\lambda_{deb}$ respectively to train the model. Here, we study the impact of $\lambda_{cons}$ and $\lambda_{deb}$ as shown in Figure 4. As $\lambda_{cons}$ rises from 0.0 to 2.0, i.e., the attribute-supervision $\mathcal{L}_{cons}$ is introduced into AENet for the semantic prediction, $H$ increases on all datasets. The best $H$ is obtained when $\lambda_{cons}=1.0$. This demonstrates the effectiveness of semantic consistency in predicting semantically relevant details, closely aligning with category prototypes to focus on attribute-specific features. When $\lambda_{cons}>1.0$, $H$ decreases. Thus, we set $\lambda_{cons}=1.0$ for optimal results. Besides, increasing $\lambda_{deb}$ leads to more consistent distribution between seen and unseen predictions, improving unseen accuracy $U$ and overall $H$ performance.
@@
To intuitively validate the capability of AENet for capturing enhanced semantic information necessary for unseen classes, we input unseen images into AENet to predict the attribute scores ($\mathcal{M}(f(x))$). As shown in Figure 5, we can observe that AENet demonstrates strong performance in predicting attributes across various animal species in the AwA2 dataset. For example, for the horse, AENet closely matches the ground truth (GT) on “brown”, “hooves”, and “domestic”, with slight discrepancies in “chewteeth” and “oldworld”. The rat predicted scores of the rat are remarkably accurate across all attributes, showcasing precision for this species. These results illustrate AENet’s robust capability in capturing enhanced semantic information across diverse species, which is crucial for generalizing to unseen classes in ZSL tasks.
@@
As shown in Figure 6, we present the t-SNE visualization of visual features for seen and unseen classes. Compared to full AENet, visual features extracted from the AENet w/o $P$ lack distinctiveness within certain classes. This intuitively suggests that the prompt plays a crucial role in enabling the application of pre-trained ViT to downstream ZSL tasks by obtaining high-quality features for seen classes. Furthermore, compared to AENet w/o $Z$, the visual features learned by our full AENet showcase desirable distinguishability with higher inter-class discrepancy and clearer decision boundaries.
@@ Conclusion
This paper introduces AENet, which leverages prompt learning and addresses the challenge of over-emphasizing primary visual features from seen domains. AENet comprises two key steps. Firstly, the concept-harmonized tokens are explored for visual and attribute modalities based on the modal-sharing token with consistent concepts. Secondly, by investigating the semantic-related details, the VRRU is proposed to distill the semantic-enhanced prompt integrated with primary visual features to attend to semantic-related information for visual enrichment. Consequently, AENet can achieve superior visual-semantic alignment for generalizing to unseen classes. Extensive experiments across three benchmark datasets show the superiority of AENet.
Acknowledgments
This work was supported in part by Fundamental Research Funds for the Central Universities (2024JBZY001, JZ2024HGTB0255, 2024XKRC082), National Natural Science Foundation of China (No. 62331003, 62120106009, 62302141, 62476021, 72434005, 92470203), Joint Funds of the National Natural Science Foundation of China under Grant U23A20314, Beijing Natural Science Foundation (L223022, L242022), Natural Science Foundation of Hebei Province (F2024105029), the Chinese Association for Artificial Intelligence (CAAI)-Compute Architecture for Neural Networks (CANN) Open Fund, developed on OpenI Community.